# Scale Reliant Inference

### v4

@@ 6 Scale Simulation Random Variables (SSRVs)
Moving beyond limits on inference, we concentrate on practical solutions to SRI. We focus our attention on Bayesian approaches as they can be applied to both Partial and Full SRI tasks. Two challenges confront the use of Bayesian PIMs for SRI: model specification and computational tractability.
We expect that, in most cases, it will be challenging to specify a likelihood model $\mathcal{L}(\theta;\mathbf{Y})$ that encodes the limited information about $\theta$ provided by $\mathbf{Y}$. For example, in general cases, we expect the estimand will not permit the same type of additive decomposition (i.e., $\theta_{d}=\theta^{\tiny\mbox{comp}}_{d}+\theta^{\tiny\mbox{tot}}$) as it does for LFC estimation (see Supplementary Section B.1 for an example related to covariance estimation). Beyond the likelihood, standard approaches to prior specification in Bayesian PIMs can imply unreasonable assumptions about the system scale. For example, for LFC estimation a prior of the form $p(\theta)$ may make implicit assumptions about $\theta^{\tiny\mbox{tot}}$ though the conditional $p(\theta^{\tiny\mbox{tot}}\mid\theta^{\tiny\mbox{comp}})$. In Section 7.2, we demonstrate how standard priors can lead to elevated rates of Type-I and Type-II errors in practice.
Even after the likelihood and prior are specified, Bayesian PIMs can be challenging to infer. Returning to LFC estimation, consider that the posterior factor $p(\theta^{\tiny\mbox{tot}}\mid\theta^{\tiny\mbox{comp}})$ retains the variance of the prior while the variance of the factor $p(\theta^{\tiny\mbox{comp}}\mid\mathbf{Y})$ will typically decrease with increasing sample size. This leads to substantial anisotropy (i.e., $\mathrm{var}(\theta^{\tiny\mbox{comp}}\mid\mathbf{Y})<<\mathrm{var}(\theta^{\tiny\mbox{tot}}\mid\mathbf{Y})$) which poses computational challenges for traditional Bayesian inferential tools (Gustafson,, 2015). While prior authors have suggested that tools like importance sampling may be used to address these problems, in our experience, those methods work poorly for SRI where posteriors are often high-dimensional.
To address these challenges, we introduce Scale Simulation Random Variables (SSRVs), a modeling framework that can help specify and infer Bayesian PIMs for SRI. We define SSRVs in the following section.
@@ 6.1 Defining SSRVs
We define SSRVs as the combination of three components. The first is the target estimand, $\theta(\mathbf{W}^{{\tiny\mbox{comp}}},W^{{\tiny\mbox{tot}}})$. This estimand also defines the problem as SII versus SRI. The second is the measurement model, $p(\mathbf{W}^{{\tiny\mbox{comp}}}\mid\mathbf{Y})$. The measurement model represents posterior uncertainty in the system composition. The third component is a scale model, $p(W^{{\tiny\mbox{tot}}}\mid\mathbf{W}^{{\tiny\mbox{comp}}})$. The scale model is a prior distribution which represents uncertainty in the system scale and may be conditioned on the system composition. With these components, a SSRV is a random variable $\theta$ with probability distribution $p(\theta\mid\mathbf{Y})$ that can be defined either algorithmically or through an integral transform.
Definition 2 (SSRVs: Algorithm).
The following algorithm simulates $S$ samples from a SSRV defined by the target estimand, measurement model, and scale model:
@@
[…]
@@
Algorithm 1 Monte-Carlo Algorithm for Simulating Scale Simulation Random Variables
Definition 3 (SSRVs: Integral Transform).
[…]
Less formally, SSRVs are simulations of the target estimand that incorporate both uncertainty in the system composition and uncertainty in the system scale. This interpretation is echoed in the SSRV mean:
@@ 6.2 Relationship to Bayesian PIMs
SSRVs are Bayesian PIMs of the form $p(\theta\mid\mathbf{Y})$ which are specified using augmented variables $W^{{\tiny\mbox{tot}}}$ and $\mathbf{W}^{{\tiny\mbox{comp}}}$. Consider that, within the integral of Equation 7, the term $p(W^{{\tiny\mbox{tot}}}\mid\mathbf{W}^{{\tiny\mbox{comp}}})p(\mathbf{W}^{{\tiny\mbox{comp}}}\mid\mathbf{Y})$ is the posterior of a Bayesian PIM. Unlike the Bayesian PIMs discussed previously, the Bayesian PIM within the integral of Equation 7 uses $W^{{\tiny\mbox{tot}}}$ and $\mathbf{W}^{{\tiny\mbox{comp}}}$ as parameters rather than $\theta$. SSRVs then use the target estimand to transform this posterior over $\mathbf{W}^{{\tiny\mbox{comp}}}$ and $W^{{\tiny\mbox{tot}}}$ into a posterior over $\theta$. SSRVs are simply an alternative method of specifying Bayesian PIMs for SRI.
@@
SSRVs mitigate the challenges of Bayesian PIMs discussed in Section 6. Using SSRVs, analysts do not need to directly specify a partially identified likelihood $\mathcal{L}(\theta;\mathbf{Y})$. Instead, analysts have the comparably easier task of specifying a fully identified likelihood $\mathcal{L}(\mathbf{W}^{{\tiny\mbox{comp}}};\mathbf{Y})$ when building the measurement model. We expect this will often be a straight-forward compositional modeling task where analysts can make use of a wide variety of available tools (e.g., Silverman et al., (2022)). Beyond the likelihood, the scale model of an SSRV provides a clear picture of modeling assumptions about scale. In Section 7.2 we highlight the benefit of this transparency using real and simulated data examples. SSRVs can also simplify posterior computations due to Algorithm 1 as the measurement model is a standard (identified) Bayesian model, and the scale model is simply a prior. Sampling from both distributions is comparatively straightforward and avoids the anisotropy common with Bayesian PIMs.
Case Study: ALDEx2 is a SSRV with a Degenerate Scale Model
@@
While there are many popular tools used for differential abundance analysis, we highlight ALDEx2 (Fernandes et al.,, 2014) as it has a particularly simple mathematical form which aligns closely with SSRVs. In viewing ALDEx2 as a SSRV, we identify the origin of its unacknowledged bias (demonstrated empirically in later sections) and identify how to improve the model and mitigate that bias.
@@
In Supplementary Section D.3, we present the ALDEx2 model as originally published in Fernandes et al., (2014). For the sake of brevity and clarity, we briefly review a simplified form of the ALDEx2 model which can be described in three steps. In the first step, posterior draws from independent multinomial-Dirichlet models are used to simulate uncertainty in the composition conditioned on the observed data. In the second step, those draws are normalized using the Centered Log-Ratio (CLR) transform which is a transform from the $D$-simplex ($\mathbb{S}^{D}$) to $\mathbb{R}^{D}$ defined as $\phi(x)=\log x-\mathbf{1}\operatorname*{mean}(\log x)$ for a vector $x\in\mathbb{S}^{D}$. In the third step, LFCs are calculated using the CLR-normalized relative abundances. The result is a distribution over LFCs. In fact, these three steps correspond to the measurement model, scale model, and target estimand of a SSRV.
@@
where $\delta[x]$ denotes a point mass centered at the point $x$. ALDEx2 does not account for any uncertainty in its implicit assumptions regarding the unmeasured system scale. Errors in this assumption lead to biased estimates.
Beyond sequence counting, we expect that unacknowledged bias due to errors in implicit assumptions regarding scale may be found in other fields. In Supplementary Section F, we provide an example of an implicit assumption in a specific type of ecological survey.
7 Empirical Studies of Sequence Counting
Through the analysis of two case-studies, we show that popular tools for differential abundance analysis display Type-I and/or Type-II error rates due to unacknowledged bias originating from problems of scale. Building on our description of ALDEx2 as a SSRV, we show that these error rates can be drastically reduced by considering error in assumptions about scale (i.e., using a scale model with non-zero variance).
@@ 7.1 ALDEx2’s Degenerate Scale Model Drives False Positives
We simulated a microbiome study comparing human gut microbiota in otherwise healthy individuals before and after treatment with a mild, narrow spectrum antibiotic whose true effect is to slightly decrease the abundance of 4 of the 20 simulated taxa (full simulation details provided in Supplementary Section G.1). Figure 1a shows the simulated true abundances ($\mathbf{W}$) for a set of $N=100$ study participants equally split between the pre- and post-antibiotic conditions. We simulated data $\mathbf{Y}$ with arbitrary scale by resampling the true abundances $\mathbf{W}$ with a multinomial distribution so the arbitrary scale of each sequenced sample $\mathbf{Y}_{\cdot n}$ has a total of $5,000$ total counts. We applied two popular DA models (DESeq2 (Love et al.,, 2014) and ALDEx2 (Fernandes et al.,, 2014)) to $\mathbf{Y}$ and, in Figure 1, show the pattern of true/false positives/negatives that result. In terms of Type-I Error control, both methods are worse than determining statistical significance by flipping a fair coin. Moreover, in Supplementary Figure 3 we show that the false discovery rate of both methods increases with increasing sample size – the hallmark of unacknowledged bias.
Figure 1: (a) Modeling results for DESeq2, ALDEx2, and a SSRV model. The Relaxed scale model is centered on the same geometric assumption as CLR normalization assumption but considers error in that assumption by using $\alpha=0.5$. b) Sensitivity analysis showing the estimated standardized log-fold change (LFC) for each taxa under the Relaxed model as a function of $\alpha$. The average and standard deviation of log fold change was computed across Monte Carlo samples, and a standardized value was computed for each entity by dividing the mean LFC by its corresponding standard deviation. A black line signifies that the estimated effect size was statistically different from zero at 95% confidence. The four taxa that were truly differentially expressed are labeled (Taxa 3, 4, 15, and 20).
@@
Our relaxation is sufficient to remove all of ALDEx2’s false positives and retain most if not all of the true positives (Figure 1a). Additionally, our relaxation addresses the unacknowledged bias and does not lead to the same asymptotic increase in false discovery rates as seen with ALDEx2 and DESeq2 (Supplemental Figure 3). Moreover, sensitivity analysis shows that our results are largely insensitive to the choice of $\alpha$: even small amounts of uncertainty are sufficient to remove the majority of false positives at the cost of just a single false negative (taxa 3; Figure 1b). Only when $\alpha$ is chosen to be implausibly large ($\alpha>1.4$) does the relaxed ALDEx2 model fail to find true positives. Considering that the CLR estimate of scale corresponds to $\hat{\theta}^{\perp}_{CLR}=0.03$ a value of $\alpha>1.4$ reflects a greater than 5% probability that the log-fold-change of scales, $\theta^{\perp}$, is outside of the range $[\exp^{-1.96\times 1.4^{2}+\hat{\theta}^{\perp}_{CLR}},\exp^{1.96\times 1.4^{2}+\hat{\theta}^{\perp}_{CLR}}]=[0.02,48.02]$; which is implausible for a narrow spectrum antibiotic. In summary, our results indicate that the CLR normalization in ALDEx2 drives its observed false positives and that these can be mitigated by considering potential errors in assumptions about scale. The observed pattern of false positives within DESeq2 suggests similar effects may be at play within that model even though it uses a different median-of-ratios normalization.
@@ 7.2 Application to a novel 16S rRNA-Seq dataset
Although artificial gut systems (bioreactors) are seeded with human feces and often used as a model for host-associated microbiota, many bioreactors undergo a two week equilibration period in which the inoculated communities diverge from those seen in human hosts and adapt to their new ex vivo environment. To evaluate the limits of these model systems, we ran and collected data from 8 bioreactors over the course of the first two weeks after inoculation of which 6 were used for analyses due to experimental constraints (see Supplemental Section G.2). Each bioreactor was seeded with 100mL of fecal slurry and maintained at a constant 400mL of growth media over the course of the experiment. Our primary interest was in determining which of the microbial taxa present in the innoculum were able to grow in the bioreactors. We quantify this interest using the differential abundance target estimator given in Equation (1) with $x_{n}=0$ denoting day 1 of the experiment and $x_{n}=1$ denoting day 14. In this case, we were primarily interested in the de facto total scale, i.e., we wanted to identify taxa that increased or decreased in total cellular number over the two week period.
@@
Five models were compared on the resultant data: one Bayesian PIM, two SSRVs, ALDEx2, and DESeq2. We used the following multinomial log-normal Bayesian PIM, which is an extension of the Bayesian multinomial logistic-normal models that have become popular for analysis of microbiota survey data (Silverman et al.,, 2022):
@@
with priors $\mathbf{B}\sim N(\mathbf{M},\boldsymbol{\Omega},\boldsymbol{\Gamma})$ and $\boldsymbol{\Omega}\sim IW(u,\boldsymbol{\Xi})$. This is a Bayesian PIM since the multinomial likelihood is only parameterized in terms of $\mathbf{W}^{{\tiny\mbox{comp}}}$ whereas prior is defined over both $\mathbf{W}^{{\tiny\mbox{comp}}}$ and $W^{{\tiny\mbox{tot}}}$. We chose the prior hyperparameters to echo standard defaults we have observed in the literature, i.e., we chose $\mathbf{M}=0$ and choose $\mathbf{\Gamma}$ and $\mathbf{\Xi}$ to be diagonal matricies (See Supplementary Section G.2 for prior specifications).
There are two critical limitations of this PIM. First, it is computationally infeasible to infer this model directly. Instead, to obtain posterior samples of the parameter $\mathbf{B}$, we had to derive an equivalent SSRV and sample $\mathbf{B}$ using Algorithm 1. Supplementary Section H derives the SSRV representation of this Bayesian PIM and shows that the measurement model is a Bayesian multinomial logistic-normal model and the scale model is a log-normal linear model. The second, more important, limitation of this model is that the assumptions about scale are not transparent. In fact, this seemingly innocuous and standard prior specification actually implies strong, biologically implausible assumptions: implicit in this prior is a non-trivial conditional relationship of the form $p(\theta^{\tiny\mbox{tot}}\mid\theta^{\tiny\mbox{comp}})$. This leads to a posterior for $\theta^{\tiny\mbox{tot}}$ which reflects a strong, biologically unreasonable belief that total microbial load is higher on day 1 than on day 14 (see Supplementary Section G.2). Below we show this leads to unacknowledged bias which presents as elevated Type-I and Type-II error rates.
@@
where $c$ is some constant and $\mu$ is subject to the constraint $\sum_{n=1}^{N}\mu_{n}=0$. Note that the LFC target estimand given by Equation 1 is invariant to the value of $c$ and so we can set $c=0$ with no loss of generality.
@@
We applied the Bayesian PIM, the two SSRVs, ALDEx2, and DESeq2 to the artificial gut data; the results are summarized in Figure 2. As before we find that ALDEx2 and DESeq2 have high inferential error rates. However, unlike the prior study where these were primarily Type-I errors, here we observe far more Type-II errors owing to the fact that most taxa are actually increasing in abundance as the community grows to occupy the larger vessels and this study has a relatively small sample size ($N=12$). Still we find the hallmark of unacknowledged bias: using resampling, Supplemental Figure 5 shows that ALDEx2 and DESeq2 display increasing rates of false discoveries with increasing sample size while consistently displaying Type-II error rates far in excess of the Design model. For example with just 20 vessels, the FDR of ALDEx2 and DESeq2 are approximately 30% and 15% respectively. In contrast, even at 50 vessels, the Design model controls FDR at a nominal 5%. As expected based on our investigation of the Bayesian PIM prior, this model displays a false discovery 45% which increases with more data (Supplemental Figure 5).
@@
Figure 2: Comparison of results from ALDEx2, DESeq2, the Bayesian PIM, and two SSRVs using real data from an ex vivo artificial gut study. “GS” denotes the gold-standard model against which true/false positives/negatives are defined. Our goal standard (the SSRV using the Flow scale model) is the only model that could use the replicate flow-cytometry measurements collected in this study. For brevity, taxa that were identified as differentially abundant with respect to any of the 5 models are not shown (based on p-value less than 0.05 or posterior 95% credible interval not covering zero).
Figure 3: Sensitivity analysis of the Design SSRV as a function of the parameter $\tau$ in the Design scale model. The mean, standard deviation, and 95% credible interval were computed using the posterior distribution of log-fold change for each entity. The standardized log-fold change (LFC) was calculated by dividing the mean log-fold change by the corresponding estimate for standard deviation. The standardized LFC as a function of $\tau$ is shown for each taxa in the artificial gut study and is colored black if the corresponding credible interval excluded zero and grey otherwise. As an illustration, we highlight the 5 taxa that are the most clearly differentially abundant in the sense that their credible sets exclude zero over the largest range of $\tau$. For comparison, we depict the value of $\tau$ used in Figure 2 ($\tau=0.5$) as a black dotted line.

### v5

@@ 3 Scale Simulation Random Variables (SSRVs)
We now discuss practical solutions to SRI. We focus our attention on Bayesian PIM approaches as they can be applied to both Partial and Full SRI tasks. Two challenges confront the use of Bayesian PIMs for SRI: model specification and computational tractability.
Standard approaches to prior specification in Bayesian PIMs can lead to unreasonable assumptions about the system scale. For LFC estimation, a prior of the form $p(\theta)$ may make implicit assumptions about $\theta^{\tiny\mbox{tot}}$ though the conditional $p(\theta^{\tiny\mbox{tot}}\mid\theta^{\tiny\mbox{comp}})$. We demonstrate this pitfall in Section 4.2.
Beyond model specification, Bayesian PIMs can be challenging to infer. Returning to LFC estimation, note the posterior factor $p(\theta^{\tiny\mbox{tot}}\mid\theta^{\tiny\mbox{comp}})$ retains the variance of the prior while the variance of the factor $p(\theta^{\tiny\mbox{comp}}\mid\mathbf{Y})$ will typically decrease with increasing sample size. This leads to substantial anisotropy (i.e., $\mathrm{var}(\theta^{\tiny\mbox{comp}}\mid\mathbf{Y})<<\mathrm{var}(\theta^{\tiny\mbox{tot}}\mid\mathbf{Y})$) which poses computational challenges for traditional Bayesian inferential tools (Gustafson, 2015). Previous authors have suggested that tools like importance sampling could address these problems (Gustafson, 2015). In our experience, those methods work poorly for SRI where posteriors are often high-dimensional.
To address these challenges, we introduce Scale Simulation Random Variables (SSRVs), a modeling framework that can help specify and infer Bayesian PIMs for SRI.
@@ 3.1 Defining SSRVs
We define SSRVs as the combination of three components. The first is the target estimand, $\theta(\mathbf{W}^{{\tiny\mbox{comp}}},W^{{\tiny\mbox{tot}}})$. This estimand also defines the problem as SII versus SRI. The second is the measurement model, $p(\mathbf{W}^{{\tiny\mbox{comp}}}\mid\mathbf{Y})$. The measurement model represents posterior uncertainty in the system composition. The third component is a scale model, $p(W^{{\tiny\mbox{tot}}}\mid\mathbf{W}^{{\tiny\mbox{comp}}})$. The scale model is a prior distribution which represents uncertainty in the system scale and may be conditioned on the system composition. With these components, a SSRV is a random variable $\theta$ with probability distribution $p(\theta\mid\mathbf{Y})$ that can be defined via an integral transform.
Definition 3.1 (SSRVs: Integral Transform).
[…]
The above integral transform suggests the following Monte-Carlo procedure to efficiently simulate SSRVs.
@@
       Sample $\mathbf{W}^{{\tiny\mbox{comp}}(s)}\sim p(\mathbf{Y})$
       Sample $W^{{\tiny\mbox{tot}}(s)}\sim p(\mathbf{W}^{{\tiny\mbox{comp}}(s)})$
       Calculate $\theta^{(s)}=\theta(\mathbf{W}^{{\tiny\mbox{comp}}(s)},W^{{\tiny\mbox{tot}}(s)})$
@@
Algorithm 1 Monte-Carlo Algorithm for Simulating SSRVs
Less formally, SSRVs are simulations of the target estimand that incorporate both uncertainty in the system composition and uncertainty in the system scale. This interpretation is echoed in the SSRV mean which is the expectation of the target estimand taken with respect to the measurement and scale models:
@@ 3.2 Relationship to Bayesian PIMs
SSRVs are Bayesian PIMs which are specified using augmented variables $W^{{\tiny\mbox{tot}}}$ and $\mathbf{W}^{{\tiny\mbox{comp}}}$. Within the integral of Equation 8, the term $p(W^{{\tiny\mbox{tot}}}\mid\mathbf{W}^{{\tiny\mbox{comp}}})p(\mathbf{W}^{{\tiny\mbox{comp}}}\mid\mathbf{Y})$ is the posterior of a Bayesian PIM. Unlike the Bayesian PIMs discussed previously, the Bayesian PIM within the integral of Equation 8 uses $W^{{\tiny\mbox{tot}}}$ and $\mathbf{W}^{{\tiny\mbox{comp}}}$ as parameters rather than $\theta$. SSRVs then use the target estimand to transform this posterior over $\mathbf{W}^{{\tiny\mbox{comp}}}$ and $W^{{\tiny\mbox{tot}}}$ into a posterior over $\theta$. SSRVs are simply an alternative method of specifying Bayesian PIMs for SRI.
@@
SSRVs mitigate the challenge of specifying and inferring Bayesian PIMs discussed above. Whereas priors placed directly on $p(\theta)$ may hide assumptions about scale within the conditional $p(\theta^{\tiny\mbox{tot}}\mid\theta^{\tiny\mbox{comp}})$, the scale model of an SSRV is that conditional. That is, assumptions about scale are an explicit part of SSRV specification. In Section 4.3 we highlight the benefit of this transparency using real and simulated data examples. SSRVs can also simplify posterior computations. Algorithm 1 only requires simulating from the measurement model and the scale model before combining those simulations in the target estimand. The measurement model is a standard (identified) Bayesian model, and the scale model is simply a prior. Sampling from both distributions is comparatively straightforward and avoids the anisotropy common with Bayesian PIMs.
4 Applications
This section considers real and simulated case studies of microbiome differential abundance analysis to illustrate the theory and methods developed in the prior sections. We start by reviewing the ALDEx2 (Fernandes et al., 2014) model which is a popular method for differential abundance analysis. We show that ALDEx2 can be thought of as an SSRV with an unrealistic scale model that underestimates prior uncertainty. Using real and simulated studies we show how this scale model leads to elevated rates of type-I and type-II errors in practice. Remarkably, we find similar patterns of false positives and false negatives with the distinct DESeq2 model (Love, Huber and Anders, 2014) suggesting our insights generalize beyond ALDEx2. We show how SSRVs address the pitfalls of these methods.
4.1 ALDEx2 is a SSRV with a Degenerate Scale Model
@@
While there are many popular tools for differential abundance analysis, we highlight ALDEx2 (Fernandes et al., 2014) as it has a particularly simple mathematical form which aligns closely with SSRVs.
@@
In Supplementary Section D.3, we present the ALDEx2 model as originally published in Fernandes et al. (2014). For brevity and clarity, we briefly review a simplified form of the ALDEx2 model which can be described in three steps. In the first step, posterior draws from independent multinomial-Dirichlet models are used to simulate uncertainty in the composition conditioned on the observed data. In the second step, those draws are normalized using the Centered Log-Ratio (CLR) transform which is a transform from the $D$-simplex ($\mathbb{S}^{D}$) to $\mathbb{R}^{D}$ defined as $\phi(x)=\log x-\mathbf{1}\operatorname*{mean}(\log x)$ for a vector $x\in\mathbb{S}^{D}$. In the third step, LFCs are calculated using the CLR-normalized relative abundances. The result is a distribution over LFCs. In fact, these three steps correspond to the measurement model, scale model, and target estimand of a SSRV.
@@
where $\delta[x]$ denotes a point mass centered at the point $x$. This scale model has zero variance. That is, ALDEx2 does not account for any uncertainty in its implicit assumption regarding the unmeasured system scale. Errors in this assumption lead to biased estimates.
As an aside, we expect that similar unacknowledged bias due to errors in implicit assumptions regarding scale may be found in other fields. Normalizations are common when analyzing polling data (Topîrceanu and Precup, 2020) and in Supplementary Section F, we provide an example of an implicit assumption in a specific type of ecological survey.
@@ 4.2 ALDEx2’s Degenerate Scale Model Drives False Positives
We simulated a microbiome study comparing human gut microbiota in otherwise healthy individuals before and after treatment with a mild, narrow spectrum antibiotic whose true effect is to slightly decrease the abundance of 4 of the 20 simulated taxa (full simulation details provided in Supplementary Section G.1). Figure 1a shows the simulated true abundances ($\mathbf{W}$) for a set of $N=100$ study participants equally split between the pre- and post-antibiotic conditions. We simulated data $\mathbf{Y}$ with arbitrary scale by resampling the true abundances $\mathbf{W}$ with a multinomial distribution so the arbitrary scale of each sequenced sample $\mathbf{Y}_{\cdot n}$ has a total of $5,000$ total counts. We applied two popular DA models (DESeq2 (Love, Huber and Anders, 2014) and ALDEx2 (Fernandes et al., 2014)) to $\mathbf{Y}$ and, in Figure 1, show the pattern of true/false positives/negatives that result. In terms of Type-I Error control, both methods are worse than determining statistical significance by flipping a fair coin. Moreover, in Supplementary Figure 3 we show that the false discovery rate of both methods increases with increasing sample size – the hallmark of unacknowledged bias.
Figure 1: (a) Modeling results for DESeq2, ALDEx2, and a SSRV model. The Relaxed scale model is centered on the same geometric assumption as CLR normalization assumption but considers error in that assumption by using $\alpha=0.5$. b) Sensitivity analysis showing the estimated standardized log-fold change (LFC) for each taxon under the Relaxed model as a function of $\alpha$. The average and standard deviation of log fold change was computed across Monte Carlo samples, and a standardized value was computed for each entity by dividing the mean LFC by its corresponding standard deviation. A black line signifies that the estimated effect size was statistically different from zero at 95% confidence. The four taxa that were truly differentially expressed are labeled (Taxa 3, 4, 15, and 20).
@@
Our relaxation is sufficient to remove all of ALDEx2’s false positives and retain most if not all of the true positives (Figure 1a). Additionally, our relaxation addresses the unacknowledged bias and does not lead to the same asymptotic increase in false discovery rates as seen with ALDEx2 and DESeq2 (Supplemental Figure 3). Moreover, sensitivity analysis shows that our results are largely insensitive to the choice of $\alpha$: even small amounts of uncertainty are sufficient to remove the majority of false positives at the cost of just a single false negative (taxon 3; Figure 1b). Only when $\alpha$ is chosen to be implausibly large ($\alpha>1.4$) does the relaxed ALDEx2 model fail to find true positives. Considering that the CLR estimate of scale corresponds to $\hat{\theta}^{\perp}_{CLR}=0.03$ a value of $\alpha>1.4$ reflects a greater than 5% probability that the log-fold-change of scales, $\theta^{\perp}$, is outside of the range $[\exp^{-1.96\times 1.4^{2}+\hat{\theta}^{\perp}_{CLR}},\exp^{1.96\times 1.4^{2}+\hat{\theta}^{\perp}_{CLR}}]=[0.02,48.02]$; which is implausible for a narrow spectrum antibiotic. In summary, our results indicate that the CLR normalization in ALDEx2 drives its observed false positives and that these can be mitigated by considering potential errors in assumptions about scale. The observed pattern of false positives within DESeq2 suggests similar effects may be at play within that model even though it uses a different median-of-ratios normalization.
@@ 4.3 Application to a novel 16S rRNA-Seq dataset
Although artificial gut systems (bioreactors) are seeded with human feces and often used as a model for host-associated microbiota, many bioreactors undergo a two week equilabration period in which the inoculated communities diverge from those seen in human hosts and adapt to their new ex vivo environment. To evaluate the limits of these model systems, we ran and collected data from 8 bioreactors over the course of the first two weeks after inoculation of which 6 were used for analyses due to experimental constraints (see Supplemental Section G.2). Each bioreactor was seeded with 100mL of fecal slurry and maintained at a constant 400mL of growth media over the course of the experiment. Our primary interest was in determining which of the microbial taxa present in the innoculum were able to grow in the bioreactors. We quantify this interest using the differential abundance target estimator given in Equation (2) with $x_{n}=0$ denoting day 1 of the experiment and $x_{n}=1$ denoting day 14. In this case, we were primarily interested in the de facto total scale, i.e., we wanted to identify taxa that increased or decreased in total cellular number over the two week period.
@@
Five models were compared on the resultant data: one specified as a Bayesian PIM, two specified as SSRVs, ALDEx2, and DESeq2. We used the following multinomial log-normal Bayesian PIM, which is an extension of the Bayesian multinomial logistic-normal models that have become popular for analysis of microbiota survey data (Silverman et al., 2022):
@@
with priors $\mathbf{B}\sim N(\mathbf{M},\boldsymbol{\Omega},\boldsymbol{\Gamma})$ and $\boldsymbol{\Omega}\sim IW(u,\boldsymbol{\Xi})$. This is a Bayesian PIM since the multinomial likelihood is only parameterized in terms of $\mathbf{W}^{{\tiny\mbox{comp}}}$ whereas prior is defined over both $\mathbf{W}^{{\tiny\mbox{comp}}}$ and $W^{{\tiny\mbox{tot}}}$. We chose the prior hyper-parameters to echo standard defaults we have observed in the literature, i.e., we chose $\mathbf{M}=0$ and choose $\mathbf{\Gamma}$ and $\mathbf{\Xi}$ to be diagonal matrices (see Supplementary Section G.2 for full prior specifications).
There are two critical limitations of this Bayesian PIM. First, it is computationally infeasible to infer this model directly. Instead, to obtain posterior samples of the parameter $\mathbf{B}$, we had to represent this model as an SSRV and sample $\mathbf{B}$ using Algorithm 1. Supplementary Section H derives the SSRV representation of this Bayesian PIM and shows that the measurement model is a Bayesian multinomial logistic-normal model and the scale model is a log-normal linear model. The second, more important, limitation of this model is that the assumptions about scale are not transparent. In fact, this seemingly innocuous and standard prior specification actually implies strong, biologically implausible assumptions: implicit in this prior is a non-trivial conditional relationship of the form $p(\theta^{\tiny\mbox{tot}}\mid\theta^{\tiny\mbox{comp}})$. This leads to a posterior for $\theta^{\tiny\mbox{tot}}$ which reflects a strong, biologically unreasonable belief that total microbial load is higher on day 1 than on day 14 (see Supplementary Section G.2). Below we show this leads to unacknowledged bias which presents as elevated Type-I and Type-II error rates.
@@
where $c$ is a constant and $\mu$ is subject to the constraint $\sum_{n=1}^{N}\mu_{n}=0$. Note that the LFC target estimand given by Equation 2 is invariant to the value of $c$ and so we can set $c=0$ with no loss of generality.
@@
We applied the Bayesian PIM, the two SSRVs, ALDEx2, and DESeq2 to the artificial gut data; the results are summarized in Figure 2. As before we find that ALDEx2 and DESeq2 have high inferential error rates. However, unlike the prior study where these were primarily Type-I errors, here we observe far more Type-II errors owing to the fact that most taxa are actually increasing in abundance as the community grows to occupy the larger vessels and this study has a relatively small sample size ($N=12$). Still we find the hallmark of unacknowledged bias: using resampling, Supplemental Figure 5 shows that ALDEx2 and DESeq2 display increasing rates of false discoveries with increasing sample size while consistently displaying Type-II error rates far in excess of the Design model. For example with just 20 vessels, the FDR of ALDEx2 and DESeq2 are approximately 30% and 15% respectively. In contrast, even at 50 vessels, the Design model controls FDR at a nominal 5%. As expected based on our investigation of the Bayesian PIM prior, the Bayesian PIM model displays a false discovery rate of 45% which further increases with more data (Supplemental Figure 5).
@@
Figure 2: Comparison of results from ALDEx2, DESeq2, the Bayesian PIM, and two SSRVs using real data from an ex vivo artificial gut study. “GS” denotes the gold-standard model against which true/false positives/negatives are defined. Our gold standard (the SSRV using the Flow scale model) is the only model that could use the replicate flow-cytometry measurements collected in this study. For brevity, taxa that were identified as differentially abundant with respect to any of the 5 models are not shown (based on p-value less than 0.05 or posterior 95% credible interval not covering zero).
Figure 3: Sensitivity analysis of the Design SSRV as a function of the parameter $\tau$ in the Design scale model. The mean, standard deviation, and 95% credible interval were computed using the posterior distribution of log-fold change for each entity. The standardized log-fold change (LFC) was calculated by dividing the mean log-fold change by the corresponding estimate for standard deviation. The standardized LFC as a function of $\tau$ is shown for each taxon in the artificial gut study and is colored black if the corresponding credible interval excluded zero and grey otherwise. As an illustration, we highlight the 5 taxa that are the most clearly differentially abundant in the sense that their credible sets exclude zero over the largest range of $\tau$. For comparison, we depict the value of $\tau$ used in Figure 2 ($\tau=0.5$) as a black dotted line.